# Stable and Accurate Orbital-Free DFT Powered by Machine Learning

### v1

@@ A.2 Model overview
As stated in the main text, we use local frames not solely to canonicalize input density coefficients with respect to global rotations, but also adopt the approach of [19] to modify the Graphormer architecture. In this modification, the known relative orientation of local frames is additionally leveraged during message passing by transforming node features from one local frame to another, allowing for the exchange of “tensorial” messages. These enable the network to effectively communicate non-scalar geometric features – something that would not be possible without the frame-to-frame transition. The representations under which internal features, i.e., the queries, keys and values in the attention mechanism, transform are hyperparameters. We choose these features to consist of 513 scalars, which remain invariant under rotations, and 85 vectors. We modify the attention mechanism of the Graphormer in the following way: when updating the features $f_{i}$ of a given node $i$, keys $k_{j}$ and values $v_{j}$ of any adjacent node $j$ are transformed into the local frame of node $i$ before computing the attention weights and aggregating the values:
$f_{i}^{(t+1)}$ $=\bigoplus_{j\in\mathcal{N}(i)}a\Big{(}q_{i},\ R(g_{i})R(g_{j}^{-1})k_{j},\ \mathbf{r}_{i}-\mathbf{r}_{j}\Big{)}\ R(g_{i})R(g_{j}^{-1})v_{j},$ (A.27)
$\mathrm{with}\quad a(q,k,\mathbf{r})$ $=\mathrm{softmax}\left(\frac{q\cdot k}{\sqrt{d}}+\mathrm{MLP}(\mathrm{GBF}(\|\mathbf{r}\|))\right),$ (A.28)
where $\mathcal{N}(i)$ is the neighborhood of node $i$. $R$ is the group representation of the keys and queries w.r.t. $\mathrm{SO}(3)$-transformations (chosen to be the same). $g_{i}$ denotes the rotation from the global frame into the local frame of node $i$. Our experiments indicate that incorporating tensorial messages, which enable direct communication of geometrical features, improves model performance (cf. Table A.1).
Table A.1: Scalar vs. tensorial messages on QM9.
@@
For our experiments on QMugs we construct the molecular graph using a radial cutoff instead of working with the fully connected graph. More specifically, we remove all edges between atoms with a pairwise Euclidean distance of $d>d_{c}$. We choose $d_{c}=6\,\mathrm{Bohr}$ as default value. The primary reason for introducing a cutoff is that message passing on the fully connected graph scales as $\mathcal{O}(N^{2})$ whereas a radial cutoff reduces the complexity to $\mathcal{O}(N)$. Additionally, we observe in our ablation study that training with cutoff leads to a better generalization of the model to larger molecules, see Table 1. Further model hyperparameters are listed in Table A.2.
Table A.2: Default model hyperparameters for our modified version of Graphormer.
@@ A.3 Enhancement modules
Natural reparametrization enables the meaningful measurement of density and gradient differences. A difference $\Delta\mathbf{p}$ in density coefficients in the LCAB ansatz (2) results in a residual density $\Delta\rho(\mathbf{r})=\sum_{\mu}\Delta p_{\mu}\omega_{\mu}(\mathbf{r})$ with an $L_{2}$ norm of
$\lVert\Delta\rho\rVert^{2}_{2}=\Delta\mathbf{p}^{\top}\mathbf{W}\Delta\mathbf{p},$ (A.29)
@@
$\lVert\Delta\rho\rVert^{2}_{2}=\lVert\Delta\tilde{\rho}\rVert^{2}_{2}=\Delta\tilde{\mathbf{p}}^{\top}\Delta\tilde{\mathbf{p}}.$ (A.30)
@@
Zhang et al. [13] propose $\mathbf{O}=\mathbf{1}$, i.e., $\mathbf{M}=\mathbf{Q}\boldsymbol{\Lambda}^{\!1\!/2}$ in their supplementary. This transformation, called canonical orthogonalization by Löwdin [61], results in highly delocalized orbitals. That is, each individual basis function aims to summarize the information in all untransformed orbitals subject to orthogonality. In our experiments, this reparametrization performed significantly worse than the symmetric version described above. This is because this type of reparametrization not only nullifies any previous transformation into local frames, but also breaks permutation invariance, rendering the energy prediction dependent on the order of atoms in the molecule. However, the code provided by [13] employs the symmetric reparametrization as we do.
The dimension-wise rescaling and atomic reference modules are implemented as in [13]. Dimension-wise rescaling linearly transforms each density coefficient independently, trading off resulting coefficient and gradient scale for each component. The atomic reference module adds a simple linear fit to the learned part of our functional. It thereby reduces the dynamic range of the predicted energy and effectively centers the gradient labels.
A.4 Model training
For both QM9 and QMugs, we train the model for 90 epochs with a batch size of 128. Over the course of all epochs, the learning rate is reduced from $7\cdot 10^{-5}$ to 00 using a cosine annealing schedule [64]. As optimizer, we use AdamW [65] with $\beta_{1}=0.95$, $\beta_{2}=0.99$ and a weight decay factor of $10^{-10}$. We do not use dropout. The training hyperparameters are summarized in Table A.3.
We apply a loss to both the energy $E$ and its gradient $\nabla_{\mathbf{p}}E$. For the energies, a simple $L_{1}$ loss is used to compare the output to the labels. The gradient needs further attention, since it is only known in the hyperplane of normalized densities. To compare the derivative of the predicted energy with the gradient label $\mathbf{g}_{\mathrm{label}}$, we follow [13] and apply an $L_{1}$ loss on the projected difference,
$\mathcal{L}_{\mathrm{gradient}}=\left\|\left(\mathbf{I}-\frac{\mathbf{w}\mathbf{w}^{T}}{\mathbf{w}^{T}\mathbf{w}}\right)\left(\mathbf{\nabla}_{\mathbf{p}}E-\mathbf{g}_{\mathrm{label}}\right)\right\|_{1},$ (A.31)
@@ A.4 Model training
Table A.3: Default training hyperparameters.
@@
A.5 Initial electron density guess: Data-driven sum of atomic densities (dSAD)
@@
A multitude of established methods for generating initial guesses for the electron density in KS-DFT exist, such as the MINAO initialization [43, 44]. However, while it is cheap compared to the Kohn-Sham iterations because of the minimal basis that it uses, it still scales cubically with system size, and additionally requires density-fitting to transform the guess to the density basis (Eq. 2).
@@
A much simpler and linear scaling option is a superposition of atomic densities (SAD). We have datasets with ground state density coefficients at hand, as these are required for training. This allows us to determine average atomic densities with a data-driven approach: We take all instances of each atom type (i.e. chemical element) in the dataset, and take the average of the corresponding coefficients over all these instances. Averages for coefficients corresponding to basis functions with $l>0$ are set to zero. For a given molecule $\mathcal{M}$, concatenating these averages for all atom types in the molecule then yields $\bar{\mathbf{p}}$, our data-driven SAD (dSAD). However, this superposition of atomic densities is not necessarily normalized to the correct electron number. Since the number of electrons stays invariant during density optimization (see section A.6), we need to normalize to generate a valid guess. One approach is to uniformly scale the coefficients linearly to the correct electron number $N_{\mathrm{e}}$, leading to
$\bar{\mathbf{p}}_{\mathrm{uniform}}=\bar{\mathbf{p}}\,\frac{N_{\mathrm{e}}}{\mathbf{w}^{\top}\bar{\mathbf{p}}}\,.$ (A.32)
@@
with the basis integrals $\mathbf{w}$, see section A.4. While simple, this has a major shortcoming: The largest part of the electron density lies close to the cores, and, for elements other than hydrogen, this core density varies only very little between different instances of the same atom type in neutral molecules. Thus, the SAD guess describes the core very precisely. The coefficients of the inner $l=0$ basis functions largely describe this core density and should hence be varied very little in the normalization. However, scaling all coefficients by the same factor to achieve normalization does not respect this. For example, the core density of atomic species with high electronegativity (whose corresponding coefficients, on average, describe a higher number of electrons than their atomic number indicates), would be scaled down and hence underestimated.
@@
This is why we propose a heteroscedastic normalization procedure, which adapts to the variance of the coefficients over the dataset: Coefficients with high variance are scaled more than those with low variance, as they are more likely to be far from the mean. As the mean $\bar{\mathbf{p}}$ and variance $\sigma_{\mu}$ of each coefficient are known, we can formulate this as an weighed least squares optimization problem with a linear constraint, where the weights of the squared deviations from $\bar{\mathbf{p}}$ are given by the inverse squares of the variances (see also Fig. A.1):
$\bar{\mathbf{p}}_{\mathrm{adaptive}}$ $=\mathop{\mathrm{arg\,max}}_{\mathbf{p},\,\,\mathbf{w}^{\top}\mathbf{p}=N_{\mathrm{e}}}\sum_{\mu}\frac{(p_{\mu}-\bar{p}_{\mu})^{2}}{2\sigma_{\mu}^{2}}=\bar{\mathbf{p}}+\mathop{\mathrm{arg\,max}}_{\mathbf{d},\,\,\mathbf{w}^{\top}\mathbf{d}=\Delta N_{\mathrm{e}}}\sum_{\mu}\frac{d_{\mu}^{2}}{2\sigma_{\mu}^{2}}$ (A.33)
@@
$\mathcal{L}(\mathbf{d},\lambda)=\sum_{\mu}\frac{d_{\mu}^{2}}{2\sigma_{\mu}^{2}}+\lambda\bigg{(}\Big{(}\sum_{\mu}w_{\mu}d_{\mu}\Big{)}-\Delta N_{\mathrm{e}}\bigg{)}\,$ (A.34)
@@
$d_{\mu}$ $=-\lambda\sigma_{\mu}^{2}w_{\mu}\,,\quad\lambda=-\frac{\Delta N_{\mathrm{e}}}{\sum_{\mu}\sigma_{\mu}^{2}w_{\mu}^{2}}\,$ (A.35)
@@
$(\bar{\mathbf{p}}_{\mathrm{adaptive}})_{\mu}=\bar{p}_{\mu}+\Delta N_{\mathrm{e}}\frac{\sigma_{\mu}^{2}w_{\mu}}{\sum_{\nu}\sigma_{\nu}^{2}w_{\nu}^{2}}\,.$ (A.36)
This result matches the intuition established above: The correction to each component of the average coefficients $\bar{\mathbf{p}}$ is proportional both to the variance of it over the dataset, and the weight of its corresponding basis function. An illustration of this method compared to simply scaling the guess is shown in Fig. A.1.
Figure A.1: Different methods for normalizing atomic densities. When extracting average atomic densities from the training set, these are not correctly normalized. Simple uniform scaling ($\bar{\mathbf{p}}_{\mathrm{uniform}}$) neglects the relative invariance of core electrons under chemical bonding. The heteroscedastic estimate $\bar{\mathbf{p}}_{\mathrm{adaptive}}$ takes the variability of different kinds of coefficients into account, see section A.5.
One could either apply this normalization per molecule, or per chemical element (in the latter case, $N_{\mathrm{e}}$ denotes the number of electrons corresponding to the atom type). We choose the latter for simplicity, normalizing the electron number per element. The guess is computed before applying natural reparametrization. Comparing our dSAD guess to the MINAO guess on 1000 QM9 molecules in Table A.4, we find that it produces guesses which are slightly closer to the ground state.
Table A.4: Density errors of initial guesses. We compare the accuracy of the MINAO [43, 44], Hückel [66] and the proposed dSAD guess (section A.5). Shown is the mean of the $L_{2}$ density error to the ground state across 1000 molecules from the QM9 dataset.
@@ A.6 Density optimization
After initialization of the electron density using our dSAD guess (section A.5), the learned energy functional enables its iterative optimization in order to find the ground state density and the corresponding energy. To conserve the number of electrons, we must not diverge from the hyperplane of normalized densities $\{\mathbf{p}\colon\mathbf{w}^{\top}\mathbf{p}=N_{e}\}$. We thus project the step $\mathbf{u}$ of the optimizer onto the hyperplane such that the density coefficients are updated according to
$\mathbf{p}^{t+1}=\mathbf{p}^{t}+\left(\mathbf{I}-\frac{\mathbf{w}\mathbf{w}^{T}}{\mathbf{w}^{T}\mathbf{w}}\right)\mathbf{u}.$ (A.37)
@@
We use gradient descent with momentum as our optimizer, with a learning rate of $0.003$ and a momentum of $0.9$ for QM9. For the QMugs dataset, we reduce the learning rate to $0.0015$. These parameters were tuned such that the density optimization shows a fast and robust convergence. The number of iterations is limited to 5000. The model is never trained on molecules from the test set, and density optimization is performed on the test set only. As Fig. 2B in the main text illustrates, density optimization on the STRUCTURES25 functional converges to gradient norms of $10^{-13}$ Ha for smaller molecules. This level of convergence is more precise than our labels are, due to imperfect density fitting and a finite convergence tolerance in the Kohn Sham calculations of our ground truth, see section A.1.2. We thus stop density optimization when the gradient norm for an entire molecule falls below $10^{-4}$ Ha.
@@
Regarding the definition of chemical accuracy, an error of 1 kcal mol-1 is a widely used threshold for acceptable energy errors. As there is no widely used definition for chemical accuracy of an electron density, we propose to compare the densities produced by different XC functionals using the same (def2-TZVP) basis set [67]. At the local density approximation (LDA) rung we use BN05[68]. Generalized gradient approximations (GGA) are represented by PBE and B97[69]. As Meta-GGA, incorporating the kinetic energy density, we use R2SCAN[24]. The highest accuracy levels are achieved by the hybrid-GGAs PBE0[70], B3LYP[71] or a meta-hybrid-GGA with PW86 exchange and B95 correlation [72, 73]. The ground state densities of 1000 randomly sampled molecules from QM9 were computed using our default Kohn Sham settings for each of the above functionals. We then measured the density difference between the XC functional PBE[51], which was used for data generation, and this assortment of functionals, with results shown in Table A.5. The density difference between PBE and methods at least at the GGA level is in the range of $7.2\cdot 10^{-4}$ to $1.3\cdot 10^{-3}$ electrons in the L2 norm, while the density difference to the less accurate LDA BN05 is around $4.1\cdot 10^{-3}$ electrons. Given the above deviations from PBE densities, and given that the PBE functional has been used in more than 200 000 publications to date, we here define “chemically accurate densities” at the level of PBE as all those that come within $7.2\cdot 10^{-4}$ electrons in the L2 norm on QM9-sized molecules.
@@
Table A.5: Density differences between PBE[51] and other XC functionals. The $L_{2}$ norm of the ground state density difference is evaluated on 1000 molecules randomly sampled from QM9.
@@
Appendix S Supplementary Text
S.1 QMugs trifluoromethoxy outliers
@@
S.2 Negative densities
The representation of the density given in Eq. 2 in principle allows for regions of negative densities to occur. When using a trained model for density optimization this could lead to problems. Since negative densities are not physical and therefore no training data with negative densities exists, it is unclear if the model can make meaningful predictions for such non-physical densities.
In practice, we report that negative densities are no failure case for our models trained on the $E_{TXC}$ target. When starting from a reasonable initial guess such as dSAD (section A.5), the negative regions of the converged densities are insignificant. This is illustrated for our QM9 model and that of M-OFDFT [13] in Fig. S.2. If negative densities were to become a problem in the future, e.g. for larger molecules or when considering more elements, one could penalize negative densities directly in the density optimization. A penalization term of the form

### v2

@@ Acknowledgments
As stated in the main text, we use local frames not solely to canonicalize input density coefficients with respect to global rotations, but also adopt a recently introduced approach [31] to modify the Graphormer architecture. In this modification, the known relative orientation of local frames is additionally leveraged during message passing by transforming node features from one local frame to another, allowing for the exchange of “tensorial” messages. These enable the network to effectively communicate non-scalar geometric features – something that would not be possible without the frame-to-frame transition. The representations under which internal features, i.e., the queries, keys and values in the attention mechanism, transform are hyperparameters. We choose these features to consist of 513 scalars, which remain invariant under rotations, and 85 vectors. We modify the attention mechanism of the Graphormer in the following way: when updating the features $f_{i}$ of a given node $i$, keys $k_{j}$ and values $v_{j}$ of any adjacent node $j$ are transformed into the local frame of node $i$ before computing the attention weights and aggregating the values:
$f_{i}^{(t+1)}$ $=\bigoplus_{j\in\mathcal{N}(i)}a\Big{(}q_{i},\ R(g_{i})R(g_{j}^{-1})k_{j},\ \mathbf{r}_{i}-\mathbf{r}_{j}\Big{)}\ R(g_{i})R(g_{j}^{-1})v_{j},$ (S.27)
$\mathrm{with}\quad a(q,k,\mathbf{r})$ $=\mathrm{softmax}\left(\frac{q\cdot k}{\sqrt{d}}+\mathrm{MLP}(\mathrm{GBF}(\|\mathbf{r}\|))\right),$ (S.28)
where $\mathcal{N}(i)$ is the neighborhood of node $i$. $R$ is the group representation of the keys and queries w.r.t. $\mathrm{SO}(3)$-transformations (chosen to be the same). $g_{i}$ denotes the rotation from the global frame into the local frame of node $i$. Our experiments indicate that incorporating tensorial messages, which enable direct communication of geometrical features, improves model performance (cf. Table S.1).
Table S.1: Scalar vs. tensorial messages on QM9.
@@
For our experiments on QMugs we construct the molecular graph using a radial cutoff instead of working with the fully connected graph. More specifically, we remove all edges between atoms with a pairwise Euclidean distance of $d>d_{c}$. We choose $d_{c}=6\,\mathrm{Bohr}$ as default value. The primary reason for introducing a cutoff is that message passing on the fully connected graph scales as $\mathcal{O}(N^{2})$ whereas a radial cutoff reduces the complexity to $\mathcal{O}(N)$. Additionally, we observe in our ablation study that training with cutoff leads to a better generalization of the model to larger molecules, see Table 1 in the main text. Further model hyperparameters are listed in Table S.2.
Table S.2: Default model hyperparameters for our modified version of Graphormer.
@@
Natural reparametrization enables the meaningful measurement of density and gradient differences. A difference $\Delta\mathbf{p}$ in density coefficients in the LCAB ansatz (Eq. 2 in the main text) results in a residual density $\Delta\rho(\mathbf{r})=\sum_{\mu}\Delta p_{\mu}\omega_{\mu}(\mathbf{r})$ with an $L_{2}$ norm of
$\lVert\Delta\rho\rVert^{2}_{2}=\Delta\mathbf{p}^{\top}\mathbf{W}\Delta\mathbf{p},$ (S.29)
@@
$\lVert\Delta\rho\rVert^{2}_{2}=\lVert\Delta\tilde{\rho}\rVert^{2}_{2}=\Delta\tilde{\mathbf{p}}^{\top}\Delta\tilde{\mathbf{p}}.$ (S.30)
@@
Zhang et al. [26] propose $\mathbf{O}=\mathbf{1}$, i.e., $\mathbf{M}=\mathbf{Q}\boldsymbol{\Lambda}^{\!1\!/2}$ in their supplementary. This transformation, called canonical orthogonalization by Löwdin [76], results in highly delocalized orbitals. That is, each individual basis function aims to summarize the information in all untransformed orbitals subject to orthogonality. In our experiments, this reparametrization performed significantly worse than the symmetric version described above. This is because this type of reparametrization not only nullifies any previous transformation into local frames, but also breaks permutation invariance, rendering the energy prediction dependent on the order of atoms in the molecule. However, the code provided by M-OFDFT [26] employs the symmetric reparametrization as we do.
The dimension-wise rescaling and atomic reference modules are implemented as in prior work [26]. Dimension-wise rescaling linearly transforms each density coefficient independently, trading off resulting coefficient and gradient scale for each component. The atomic reference module adds a simple linear fit to the learned part of our functional. It thereby reduces the dynamic range of the predicted energy and effectively centers the gradient labels.
S.4 Model training
For both QM9 and QMugs, we train the model for 90 epochs with a batch size of 128. Over the course of all epochs, the learning rate is reduced from $7\cdot 10^{-5}$ to 00 using a cosine annealing schedule [79]. As optimizer, we use AdamW [80] with $\beta_{1}=0.95$, $\beta_{2}=0.99$ and a weight decay factor of $10^{-10}$. We do not use dropout. The training hyperparameters are summarized in Table S.3.
We apply a loss to both the energy $E$ and its gradient $\nabla_{\mathbf{p}}E$. For the energies, a simple $L_{1}$ loss is used to compare the output to the labels. The gradient needs further attention, since it is only known in the hyperplane of normalized densities. To compare the derivative of the predicted energy with the gradient label $\mathbf{g}_{\mathrm{label}}$, we follow M-OFDFT [26] and apply an $L_{1}$ loss on the projected difference,
$\mathcal{L}_{\mathrm{gradient}}=\left\|\left(\mathbf{I}-\frac{\mathbf{w}\mathbf{w}^{T}}{\mathbf{w}^{T}\mathbf{w}}\right)\left(\mathbf{\nabla}_{\mathbf{p}}E-\mathbf{g}_{\mathrm{label}}\right)\right\|_{1},$ (S.31)
@@
Table S.3: Default training hyperparameters.
@@
S.5 Initial electron density guess: Data-driven sum of atomic densities (dSAD)
@@
A multitude of established methods for generating initial guesses for the electron density in KS-DFT exist, such as the MINAO initialization [55, 56]. However, while it is cheap compared to the Kohn-Sham iterations because of the minimal basis that it uses, it still scales cubically with system size, and additionally requires density-fitting to transform the guess to the density basis (Eq. 2 in the main text).
@@
A much simpler and linear scaling option is a superposition of atomic densities (SAD). We have datasets with ground state density coefficients at hand, as these are required for training. This allows us to determine average atomic densities with a data-driven approach: We take all instances of each atom type (i.e. chemical element) in the dataset, and take the average of the corresponding coefficients over all these instances. Averages for coefficients corresponding to basis functions with $l>0$ are set to zero. For a given molecule $\mathcal{M}$, concatenating these averages for all atom types in the molecule then yields $\bar{\mathbf{p}}$, our data-driven SAD (dSAD). However, this superposition of atomic densities is not necessarily normalized to the correct electron number. Since the number of electrons stays invariant during density optimization (see section S.6), we need to normalize to generate a valid guess. One approach is to uniformly scale the coefficients linearly to the correct electron number $N_{\mathrm{e}}$, leading to
$\bar{\mathbf{p}}_{\mathrm{uniform}}=\bar{\mathbf{p}}\,\frac{N_{\mathrm{e}}}{\mathbf{w}^{\top}\bar{\mathbf{p}}}\,.$ (S.32)
@@
with the basis integrals $\mathbf{w}$, see section S.4. While simple, this has a major shortcoming: The largest part of the electron density lies close to the cores, and, for elements other than hydrogen, this core density varies only very little between different instances of the same atom type in neutral molecules. Thus, the SAD guess describes the core very precisely. The coefficients of the inner $l=0$ basis functions largely describe this core density and should hence be varied very little in the normalization. However, scaling all coefficients by the same factor to achieve normalization does not respect this. For example, the core density of atomic species with high electronegativity (whose corresponding coefficients, on average, describe a higher number of electrons than their atomic number indicates), would be scaled down and hence underestimated.
@@
This is why we propose a heteroscedastic normalization procedure, which adapts to the variance of the coefficients over the dataset: Coefficients with high variance are scaled more than those with low variance, as they are more likely to be far from the mean. As the mean $\bar{\mathbf{p}}$ and variance $\sigma_{\mu}$ of each coefficient are known, we can formulate this as an weighed least squares optimization problem with a linear constraint, where the weights of the squared deviations from $\bar{\mathbf{p}}$ are given by the inverse squares of the variances (see also Fig. S.1):
$\bar{\mathbf{p}}_{\mathrm{adaptive}}$ $=\mathop{\mathrm{arg\,max}}_{\mathbf{p},\,\,\mathbf{w}^{\top}\mathbf{p}=N_{\mathrm{e}}}\sum_{\mu}\frac{(p_{\mu}-\bar{p}_{\mu})^{2}}{2\sigma_{\mu}^{2}}=\bar{\mathbf{p}}+\mathop{\mathrm{arg\,max}}_{\mathbf{d},\,\,\mathbf{w}^{\top}\mathbf{d}=\Delta N_{\mathrm{e}}}\sum_{\mu}\frac{d_{\mu}^{2}}{2\sigma_{\mu}^{2}}$ (S.33)
@@
$\mathcal{L}(\mathbf{d},\lambda)=\sum_{\mu}\frac{d_{\mu}^{2}}{2\sigma_{\mu}^{2}}+\lambda\bigg{(}\Big{(}\sum_{\mu}w_{\mu}d_{\mu}\Big{)}-\Delta N_{\mathrm{e}}\bigg{)}\,$ (S.34)
@@
$d_{\mu}$ $=-\lambda\sigma_{\mu}^{2}w_{\mu}\,,\quad\lambda=-\frac{\Delta N_{\mathrm{e}}}{\sum_{\mu}\sigma_{\mu}^{2}w_{\mu}^{2}}\,$ (S.35)
@@
$(\bar{\mathbf{p}}_{\mathrm{adaptive}})_{\mu}=\bar{p}_{\mu}+\Delta N_{\mathrm{e}}\frac{\sigma_{\mu}^{2}w_{\mu}}{\sum_{\nu}\sigma_{\nu}^{2}w_{\nu}^{2}}\,.$ (S.36)
This result matches the intuition established above: The correction to each component of the average coefficients $\bar{\mathbf{p}}$ is proportional both to the variance of it over the dataset, and the weight of its corresponding basis function. An illustration of this method compared to simply scaling the guess is shown in Fig. S.1.
Figure S.1: Different methods for normalizing atomic densities. When extracting average atomic densities from the training set, these are not correctly normalized. Simple uniform scaling ($\bar{\mathbf{p}}_{\mathrm{uniform}}$) neglects the relative invariance of core electrons under chemical bonding. The heteroscedastic estimate $\bar{\mathbf{p}}_{\mathrm{adaptive}}$ takes the variability of different kinds of coefficients into account, see section S.5.
One could either apply this normalization per molecule, or per chemical element (in the latter case, $N_{\mathrm{e}}$ denotes the number of electrons corresponding to the atom type). We choose the latter for simplicity, normalizing the electron number per element. The guess is computed before applying natural reparametrization. Comparing our dSAD guess to the MINAO guess on 1000 QM9 molecules in Table S.4, we find that it produces guesses which are slightly closer to the ground state.
Table S.4: Density errors of initial guesses. We compare the accuracy of the MINAO [55, 56], Hückel [81] and the proposed dSAD guess (section S.5). Shown is the mean of the $L_{2}$ density error to the ground state across 1000 molecules from the QM9 dataset.
@@
After initialization of the electron density using our dSAD guess (section S.5), the learned energy functional enables its iterative optimization in order to find the ground state density and the corresponding energy. To conserve the number of electrons, we must not diverge from the hyperplane of normalized densities $\{\mathbf{p}\colon\mathbf{w}^{\top}\mathbf{p}=N_{e}\}$. We thus project the step $\mathbf{u}$ of the optimizer onto the hyperplane such that the density coefficients are updated according to
$\mathbf{p}^{t+1}=\mathbf{p}^{t}+\left(\mathbf{I}-\frac{\mathbf{w}\mathbf{w}^{T}}{\mathbf{w}^{T}\mathbf{w}}\right)\mathbf{u}.$ (S.37)
We use gradient descent with momentum as our optimizer, with a learning rate of $0.003$ and a momentum of $0.9$ for QM9. For the QMugs dataset, we reduce the learning rate to $0.0015$. These parameters were tuned such that the density optimization shows a fast and robust convergence. The number of iterations is limited to 5000.
Note that while KS-DFT typically requires significantly fewer steps to converge, it relies on much more complex SCF iterations. While OF-DFT requires more steps, each is computationally inexpensive. This results in a significant net gain in efficiency, particularly for larger systems, see section S.11 for details.
@@
The model is never trained on molecules from the test set, and density optimization is performed on the test set only. As Fig. 2B in the main text illustrates, density optimization on the STRUCTURES25 functional converges to gradient norms of $10^{-13}$ Ha for smaller molecules. This level of convergence is more precise than our labels are, due to imperfect density fitting and a finite convergence tolerance in the Kohn Sham calculations of our ground truth, see section S.1.2. We thus stop density optimization when the gradient norm for an entire molecule falls below $10^{-4}$ Ha.
@@
Regarding the definition of chemical accuracy, an error of 1 kcal mol-1 is a widely used threshold for acceptable energy errors. As there is no widely used definition for chemical accuracy of an electron density, we propose to compare the densities produced by different XC functionals using the same (def2-TZVP) basis set [82]. At the local density approximation (LDA) rung we use BN05[83]. Generalized gradient approximations (GGA) are represented by PBE and B97[84]. As Meta-GGA, incorporating the kinetic energy density, we use R2SCAN[36]. The highest accuracy levels are achieved by the hybrid-GGAs PBE0[85], B3LYP[86] or a meta-hybrid-GGA with PW86 exchange and B95 correlation [87, 88]. The ground state densities of 1000 randomly sampled molecules from QM9 were computed using our default Kohn Sham settings for each of the above functionals. We then measured the density difference between the XC functional PBE[66], which was used for data generation, and this assortment of functionals, with results shown in Table S.5. The density difference between PBE and methods at least at the GGA level is in the range of $7.2\cdot 10^{-4}$ to $1.3\cdot 10^{-3}$ electrons in the L2 norm, while the density difference to the less accurate LDA BN05 is around $4.1\cdot 10^{-3}$ electrons. Given the above deviations from PBE densities, and given that the PBE functional has been used in more than 200 000 publications to date, we here define “chemically accurate densities” at the level of PBE as all those that come within $7.2\cdot 10^{-4}$ electrons in the L2 norm on QM9-sized molecules.
@@
Table S.5: Density differences between PBE[66] and other XC functionals. The $L_{2}$ norm of the ground state density difference is evaluated on 1000 molecules randomly sampled from QM9.
@@
Figure S.2: The residual between STRUCTURES25 and its OF training labels is smaller than the error incurred by density fitting training labels to the KS calculations. The sketch is to scale and illustrates the pairwise $L_{2}$ norm per electron between ground state densities, averaged over the QM9 test set.
S.7 QMugs trifluoromethoxy outliers
@@
To further analyze the phenomenon, we evaluated the accuracy of STRUCTURES25 on smaller molecules containing a trifluoromethoxy group. These exhibited large errors as well, see Table S.6. Finally, the density error for a larger molecule containing the group is visualized in Figure S.4, demonstrating that the error is spatially localized around the group in question. Taken together, these observations support the hypothesis that the lack of sufficiently many training molecules including the trifluoromethoxy group is the root cause of this failure mode. Analyses of this type can help identify and close gaps in the training data.
Table S.6: Energy error of STRUCTURES25 relative to PBE/6-31G(2df,p) for small trifluoromethoxy molecules and their substitues. For the substitues, the trifluoromethoxy moiety was replaced with a methoxy group.
Molecule $\Delta E$ [mHa] substitute $\Delta E$ [mHa]
Trifluoro(methoxy)methane 629.21 1.57
(Trifluoromethoxy)benzene 642.50 1.26
2-(Trifluoromethoxy)aniline 639.28 1.81
Figure S.4: For one of the energy prediction outliers from Fig. 4 in the main text: The density prediction error is fully localized around the trifluoromethoxy group. Isocontours of the density residual between STRUCTURES25 prediction and ground truth after density fitting. The red/blue surfaces show positive/negative deviations above $0.02\,\mathrm{Bohr^{-3}}$ and the discrepancy is localized on the trifluoromethoxy group on the right.
S.8 Negative densities
The representation of the density given in Eq. 2 in the main text in principle allows for regions of negative densities to occur. When using a trained model for density optimization this could lead to problems. Since negative densities are not physical and therefore no training data with negative densities exists, it is unclear if the model can make meaningful predictions for such non-physical densities.
In practice, we report that negative densities are no failure case for our models trained on the $E_{TXC}$ target. When starting from a reasonable initial guess such as dSAD (section S.5), the negative regions of the converged densities are insignificant. This is illustrated for our QM9 model and that of M-OFDFT [26] in Fig. S.5. If negative densities were to become a problem in the future, e.g. for larger molecules or when considering more elements, one could penalize negative densities directly in the density optimization. A penalization term of the form